### v1

@@ 4 Nonlinear Modified PageRank problem
$\sum_{\{u,v\}\in\partial(\mathcal{S}_{*})}\left(\left(x(u)-x(v)\right)^{2}+\frac{\zeta}{\left(1+\sigma\right)^{2}}\right)^{\frac{1}{2}}<\left(\frac{1}{1+\sigma}\right)|\partial(\mathcal{C})|.$
Since $x(u)-x(v)=0$ cannot be discounted for any $v\sim u$ with $u,v\in\mathcal{V}\setminus\{s\}$, it follows that
@@
$\sum_{\{u,v\}\in\partial(\mathcal{S}_{*})}\left(\left(x(u)-x(v)\right)^{2}+\frac{\zeta}{\left(1+\sigma\right)^{2}}\right)^{\frac{1}{2}}\geq\tau|\partial(\mathcal{S}_{*})|.$
@@
$\tau|\partial(\mathcal{S}_{*})|<\left(\frac{1}{1+\sigma}\right)|\partial(\mathcal{C})|.$
A greater value of $\epsilon$ has the effect of increasing the value of $\sigma$, which in turn produces a tighter bound. It is worth keeping in mind that this result only applies in a perturbation context.
@@ 5 Numerical methodology
Consider $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ with $\mathcal{E}\neq\emptyset$. Again, $n=|\mathcal{V}|$ and $m=|\mathcal{E}|$. Let $g(x)=\beta r-Tf(x)$. For the purpose of obtaining a numerical solution, the Nonlinear Modified PageRank problem is re-expressed as follows: Given $g:\mathbb{R}^{n}\to\mathbb{R}^{n}$, find $x\in\mathbb{R}^{n}$ that satisfies
(11) $g(x)=0.$
@@
(12) $\min_{x\in\mathbb{R}^{n}}\psi(x)$
solves for $x^{*}$. $\psi(x)$ is referred to as a (least-squares) merit function.
Remark 5.1.
Even though a local minimiser of $\psi$ need not be a solution of (11), the merit function $\psi$ has nonetheless been used successfully in practice [26, Chapter 11.2].
@@ 5.1 Levenberg-Marquardt method
Optimisation algorithms proceed from an initial iterate $x_{0}$ and generate a finite sequence of iterates $\{x_{k}\}_{k>0}$ that either converge towards a sufficiently accurate approximation of the solution $x^{*}$, or represent the failed attempt to do so. What distinguishes one algorithm from another is the transition from a current iterate $x_{c}$ to the next iterate $x_{+}$.
@@
for $j,k=1,2,\ldots,n$. Summation is implied over the repeated index.
Corollary 5.4.
The Jacobian $J$ is Lipschitz continuous.
Proof 5.5.
Let $y\in\mathbb{R}^{n}$ and $\mu\in\mathbb{R}$. Define $a=y+\mu(x-y)$. Then
$\frac{\partial}{\partial\mu}J(y+\mu(x-y))=\frac{\partial J}{\partial a}\frac{\partial a}{\partial\mu}=J^{\prime}(y+\mu(x-y))\cdot(x-y).$
@@
[…]
Take $\mathtt{\mathbf{J}}$ to be a matricization of the third-order tensor $J^{\prime}(y+\mu(x-y))$. Note that
$\lVert J^{\prime}(y+\mu(x-y))\cdot(x-y)\rVert_{2}\leq\lVert J^{\prime}(y+\mu(x-y))\cdot(x-y)\rVert_{\text{F}}=\lVert\mathtt{\mathbf{J}}(x-y)\rVert_{2}$
and $\lVert\mathtt{\mathbf{J}}(x-y)\rVert_{2}\leq\lVert\mathtt{\mathbf{J}}\rVert_{\text{F}}\lVert x-y\rVert_{2}=\lVert J^{\prime}(y+\mu(x-y))\rVert_{\text{F}}\lVert x-y\rVert_{2}$. Subsequently,
[…]
$\leq\xi\left\lVert x-y\right\rVert_{2},$
where $\xi=\max_{b\in[x,y]}\lVert J^{\prime}(b)\rVert_{\text{F}}$. This satisfies the Lipschitz condition.
Corollary 5.6.
@@
Next, $\text{ker}\,(B)=\text{ker}\,(B^{\dagger}KB)$ is tackled. If $x\in\text{ker}\,(B)$, then $Bx=0$. Thus, $B^{\dagger}KBx=0$, and $x\in\text{ker}\,(B^{\dagger}KB)$. Conversely, if $x\in\text{ker}\,(B^{\dagger}KB)$, then $B^{\dagger}KBx=0$. The null space of $B^{\dagger}$ coincides with that of $B^{\top}$. It therefore suffices to consider $B^{\top}KBx=0$. Take $\kappa$ to be the smallest of all the (positive) diagonal entries of $K$. Subsequently, $\kappa(Bx)^{\top}(Bx)\leq(Bx)^{\top}K(Bx)=0$. Thus, $Bx=0$, and $x\in\text{ker}\,(B)$.
Finally, $\text{ker}\,(B)=\text{ker}\,(TB^{\dagger}KB)$ can be shown. If $x\in\text{ker}\,(B)$, then $Bx=0$. Thus, $TB^{\dagger}KBx=0$, and $x\in\text{ker}\,(TB^{\dagger}KB)$. Conversely, if $x\in\text{ker}\,(TB^{\dagger}KB)$, then $TB^{\dagger}KBx=0$. But $T^{-1}TB^{\dagger}KBx=B^{\dagger}KBx=0$. This gives $Bx=0$, and $x\in\text{ker}\,(B)$. By Proposition 2.2, the Jacobian is of rank $n-1$.
@@
    Thus, $y^{\top}Hy>0$.
A necessary condition (refer to [26, Theorem 2.2] and [18, Theorem 1.3.1]) for a local minimiser $x_{t}$ of $\theta_{c}$ is
@@
(13) $x_{t}=x_{c}-\left(J(x_{c})^{\top}J(x_{c})+\lambda_{c}I\right)^{-1}J(x_{c})^{\top}g(x_{c}).$
$x_{t}$ is treated as a trial solution that could be the next iterate $x_{+}$ depending on how well the quadratic model approximates $\psi$. This is gauged through a comparison of the actual reduction
@@
If $\varrho=1$, then the quadratic model faithfully reproduces the behaviour of $\psi(x)$ around $x_{c}$. The trial solution is accepted for a value of $\varrho$ sufficiently greater than zero, and $\lambda_{c}$ is possibly decreased. Otherwise, $\lambda_{c}$ is increased and a new trial solution is computed. An algorithm detailing the steps in each iteration of the Levenberg-Marquardt method can be found in [18, Algorithm 3.3.5]. Quadratic convergence can be experienced [18, Theorem 3.3.4]. However, this is contingent on the sequence $\{x_{k}\}$ having the limit $x^{*}$ and $J(x^{*})$ being full rank.
Remark 5.10.
@@ 5.2 Rank deficiency and the Jacobian
Ipsen et al. looked into the issue of a rank-deficient Jacobian in nonlinear least-squares problems [17] but for small $\psi(x^{*})$. They recommend subset selection applied to the Jacobian over the truncated singular value decomposition. This involves the formation of a full rank variant, say $\widetilde{J}$, comprised solely of linearly independent columns from $J$. Instead of relying on (13), the system of linear equations
@@
There is the outstanding question though of how to discover the linearly independent columns of the Jacobian in the first place. A rank revealing algorithm could be employed for this task, but there is no need. After all, the underlying graph-based nature of the Nonlinear Modified PageRank problem can be leveraged to determine the column in the Jacobian (see Corollary 5.6) that will play no part in the full rank variant. Since the problem solution figures in the graph partitioning process, it seems most prudent to fix the value of $x(v)$ for $v\in\mathcal{V}$ that is the furthest (minimum weighted) distance from the starting vertex $s\in\mathcal{V}$. Clearly, the relevant column in the Jacobian would then be the one omitted from $\widetilde{J}$.
Remark 5.11.
An alternative strategy could be to exclude the column in the Jacobian corresponding to the smallest entry of the minimal least squares solution to the Nonlinear Modified PageRank problem at $p=2$, i.e.,
$x=\beta Y^{\dagger}r,$
where $Y=TB^{\dagger}B$.

### v2

@@ 4 Nonlinear PageRank problem
$\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}=(1+\sigma)q+\delta,$
where
$\sigma=\frac{1}{|\mathcal{E}|}\sum_{i=1}^{|\mathcal{E}|}\left(\left(Bx^{0}\right)_{i}^{2}+\zeta\right)^{-\frac{\epsilon}{2}}-1$
and $\delta$ is the deviation.
Denote the sweep set based on the Perturbed Nonlinear PageRank problem that returns the smallest conductance as $\mathcal{S}_{*}^{\epsilon}\subset\mathcal{V}$. The volume of subsets $\mathcal{F}$ and $\mathcal{S}_{*}^{\epsilon}$ are to be related by $\text{vol}\,(\mathcal{F})=\tau\text{vol}\,(\mathcal{S}_{*}^{\epsilon})$, in which $\tau\in\mathbb{R}_{>0}$. Let $\zeta=\varepsilon\left(1+\sigma\right)^{2}\tau$ with $\varepsilon\in\mathbb{R}_{>0}$. For a sufficiently small $\lVert\delta\rVert$ along with $\text{vol}(\mathcal{F})\leq\text{vol}(\mathcal{V}\setminus\mathcal{F})$ and $\text{vol}\,(\mathcal{S}_{*}^{\epsilon})\leq\text{vol}\,(\mathcal{V}\setminus\mathcal{S}_{*}^{\epsilon})$,
$\Phi(\mathcal{S}_{*}^{\epsilon})<\frac{1}{\varepsilon(1+\sigma)^{2}}\Phi(\mathcal{F}).$
Proof.
First of all, note that $\sigma$ represents a small non-negative value. Based on Proposition 4.3, $Bx^{0}=Bc$. Since $T^{-1}\geqslant 0$, as stated in Proposition 3.1, and $r(v)\geq 0$ for all $v\in\mathcal{V}$, $c(v)\geq 0$ for all $v\in\mathcal{V}$. Furthermore, $c(v)<1$ for all $v\in\mathcal{V}$ according to Proposition 3.2, and thus the entries of $|Bx^{0}|$ are less than $1$. The assumption that $(Bx^{0})(v)+\zeta\leq 1$ for all $v\in\mathcal{V}$ together with $\epsilon\ll 1$ facilitates a small non-negative $\sigma$.
For $\delta=0$, the Perturbed Nonlinear PageRank problem (5) has the form
$(1+\sigma)TB^{\dagger}By(\epsilon)=\beta r.$
The least squares solution $(1+\sigma)y(\epsilon)=c-(1/n)q$ follows from Proposition 4.3 and
$y(\epsilon)=\frac{1}{1+\sigma}c-\frac{1}{n(1+\sigma)}q.$
Let $M=\text{diag}\,(\delta_{1},\ldots,\delta_{|\mathcal{E}|})$ and $N=(1+\sigma)B^{\dagger}B+B^{\dagger}MB$. Then, the Perturbed Nonlinear PageRank problem can be written as $TNx(\epsilon)=\beta r$. Suppose that $\kappa q+z\in\text{ker}\,(N)$ with $q^{\top}z=0$, where $\kappa\in\mathbb{R}$ and $z\in\mathbb{R}^{n}$, is an orthogonal decomposition. However,
$0=N(\kappa q+z)=\kappa Nq+Nz=\kappa(1+\sigma)B^{\dagger}Bq+\kappa B^{\dagger}MBq+Nz=Nz,$
as $Bq=0$. Since $B^{\dagger}B=I-(1/n)qq^{\top}$ from Proposition 4.2 and $q^{\top}z=0$, it follows that
$0=Nz=(1+\sigma)z+B^{\dagger}MBz.$
Hence,
$\lVert z\rVert=\frac{1}{1+\sigma}\left\lVert B^{\dagger}MBz\right\rVert.$
In view of $\lVert M\rVert=\max_{1\leq i\leq|\mathcal{E}|}|\delta_{i}|\leq\lVert\delta\rVert$,
$\lVert z\rVert\leq\frac{1}{1+\sigma}\lVert B^{\dagger}\rVert\lVert\delta\rVert\lVert B\rVert\lVert z\rVert.$
A sufficiently small $\lVert\delta\rVert$, i.e.,
$\frac{1}{1+\sigma}\lVert B^{\dagger}\rVert\lVert\delta\rVert\lVert B\rVert<1,$
implies $\lVert z\rVert=0$. Thus, $\text{ker}\,(N)=\text{span}\,(q)$ and so $\text{rank}(N)=n-1$.
Observe that $\text{rank}\,(N)=\text{rank}\,(B^{\dagger}B)$. This is a consequence of $\text{ker}\,(B^{\dagger}B)=\text{ker}\,(B)$ and Proposition 2.1. If $z\in\text{ker}\,(B^{\dagger}B)$, then $B^{\dagger}Bz=0$ and $Bz\in\text{ker}\,(B^{\dagger})$. But $\text{ker}\,(B^{\dagger})=\text{ker}\,(B^{\top})$, so $B^{\top}Bz=0$. Consider $z^{\top}B^{\top}Bz=\lVert Bz\rVert^{2}=0$. Thus, $Bz=0$ and $z\in\text{ker}\,(B)$. Conversely, if $z\in\text{ker}\,(B)$, then $Bz=0$. This results in $B^{\dagger}Bz=0$ and $z\in\text{ker}\,(B^{\dagger}B)$.
The bounded difference between $x(\epsilon)$ and $y(\epsilon)$ is
$\lVert x(\epsilon)-y(\epsilon)\rVert=\left\lVert N^{\dagger}c-\left((1+\sigma)B^{\dagger}B\right)^{\dagger}c\right\rVert\leq\left\lVert N^{\dagger}-\left((1+\sigma)B^{\dagger}B\right)^{\dagger}\right\rVert\lVert c\rVert.$
Because $\lVert\delta\rVert$ is small, $N$ can be treated as a perturbed matrix. Then, by [26, Theorem 3.4 on p. 645],
$\left\lVert N^{\dagger}-\left((1+\sigma)B^{\dagger}B\right)^{\dagger}\right\rVert$ $\leq\frac{1+\sqrt{5}}{2}\left\lVert\left((1+\sigma)B^{\dagger}B\right)^{\dagger}\right\rVert\left\lVert N^{\dagger}\right\rVert\left\lVert B^{\dagger}MB\right\rVert$
$=\frac{1+\sqrt{5}}{2(1+\sigma)}\left\lVert B^{\dagger}B\right\rVert\left\lVert N^{\dagger}\right\rVert\left\lVert B^{\dagger}MB\right\rVert,$
where the resulting equality makes use of Proposition 2.3. For any $\{u,v\}\in\mathcal{E}$,
$\left|x^{\epsilon}(u)-x^{\epsilon}(v)-\left(y^{\epsilon}(u)-y^{\epsilon}(v)\right)\right|$ $\leq\left|x^{\epsilon}(u)-y^{\epsilon}(u)\right|+\left|x^{\epsilon}(v)-y^{\epsilon}(v)\right|$
$\leq 2\lVert x(\epsilon)-y(\epsilon)\rVert$
$\leq\frac{1+\sqrt{5}}{1+\sigma}\left\lVert B^{\dagger}B\right\rVert\left\lVert N^{\dagger}\right\rVert\left\lVert B^{\dagger}\right\rVert\lVert\delta\rVert\lVert B\rVert\lVert c\rVert$
$\leq\frac{1+\sqrt{5}}{1+\sigma}\left\lVert N^{\dagger}\right\rVert\left\lVert B^{\dagger}\right\rVert^{2}\lVert B\rVert^{2}\lVert\delta\rVert\lVert c\rVert,$
in which $x^{\epsilon}=x(\epsilon)$ and $y^{\epsilon}=y(\epsilon)$. Subsequently,
@@
$\left|x^{\epsilon}(u)-x^{\epsilon}(v)-\frac{c(u)-c(v)}{1+\sigma}\right|=\mathcal{O}(\lVert\delta\rVert).$
The cornerstone of the concluding argument rests upon $\sum_{\{u,v\}\in\partial(\mathcal{S}_{*}^{\epsilon})}((c(u)-c(v))^{2}+\zeta)$. An assumption of $(Bx^{0})(v)+\zeta\leq 1$ for all $v\in\mathcal{V}$ supports
$\sum_{\{u,v\}\in\partial(\mathcal{S}_{*}^{\epsilon})}\left(\left((1+\sigma)\left(x^{\epsilon}(u)-x^{\epsilon}(v)\right)+\mathcal{O}\left(\lVert\delta\rVert\right)\right)^{2}+\zeta\right)<|\partial(\mathcal{F})|.$
It follows that
$\sum_{\{u,v\}\in\partial(\mathcal{S}_{*}^{\epsilon})}\left(\left(x^{\epsilon}(u)-x^{\epsilon}(v)\right)^{2}+\mathcal{O}\left(\lVert\delta\rVert\right)+\frac{\zeta}{\left(1+\sigma\right)^{2}}\right)<\frac{1}{\left(1+\sigma\right)^{2}}|\partial(\mathcal{F})|.$
As $x^{\epsilon}(u)-x^{\epsilon}(v)=0$ cannot be discounted for any $\{u,v\}\in\partial(\mathcal{S}_{*}^{\epsilon})$,
$\sum_{\{u,v\}\in\partial(\mathcal{S}_{*}^{\epsilon})}\left(\left(x^{\epsilon}(u)-x^{\epsilon}(v)\right)^{2}+\mathcal{O}\left(\lVert\delta\rVert\right)+\frac{\zeta}{\left(1+\sigma\right)^{2}}\right)\geq\varepsilon\tau|\partial(\mathcal{S}_{*}^{\epsilon})|+\mathcal{O}\left(\lVert\delta\rVert\right),$
@@
$\varepsilon\tau|\partial(\mathcal{S}_{*}^{\epsilon})|<\frac{1}{(1+\sigma)^{2}}|\partial(\mathcal{F})|\qquad\text{and}\qquad\Phi(\mathcal{S}_{*}^{\epsilon})<\frac{1}{\varepsilon(1+\sigma)^{2}}\Phi(\mathcal{F}).\qed$
@@ 5 Numerical methodology
Consider $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ with $\mathcal{E}\neq\emptyset$. Again, $n=|\mathcal{V}|$ and $m=|\mathcal{E}|$. Let $g(x)=\beta r-Tf(x)$. For the purpose of obtaining a numerical solution, the Nonlinear PageRank problem is re-expressed as follows: Given $g:\mathbb{R}^{n}\to\mathbb{R}^{n}$, find $x\in\mathbb{R}^{n}$ that satisfies
$g(x)=0.$ (6)
@@
$\min_{x\in\mathbb{R}^{n}}\psi(x)$ (7)
solves for $x^{*}$. $\psi(x)$ is referred to as a (least squares) merit function.
Remark 4.
Even though a local minimiser of $\psi$ need not be a solution of (6), the merit function $\psi$ has nonetheless been used successfully in practice [27, Chapter 11.2].
@@ 5.1 Levenberg-Marquardt method
Optimisation algorithms proceed from an initial iterate $x_{0}$ and generate a finite sequence of iterates $\{x_{k}\}_{k>0}$ that either converge towards a sufficiently accurate approximation of the solution $x^{*}$, or represent the failed attempt to do so. Algorithms are distinguished by the transition from a current iterate $x_{c}$ to the next iterate $x_{+}$.
@@
for $j,k=1,2,\ldots,n$. Summation is implied over the repeated index. ∎
Corollary 5.1.1.
The Jacobian $J$ is (locally) Lipschitz continuous.
Proof.
Define $a=y+\mu(x-y)$, in which $y\in\mathbb{R}^{n}$ and $\mu\in[0,1]$. Then
$\frac{d}{d\mu}J(y+\mu(x-y))=\frac{\partial J}{\partial a}\frac{da}{d\mu}=J^{\prime}(a(\mu))\bar{\times}_{3}(x-y).$
@@
$J(x)-J(y)=\int_{0}^{1}\,J^{\prime}(a(\mu))\bar{\times}_{3}(x-y)\,d\mu.$
Let $\mathtt{\mathbf{J}}$ be a matricization of $J^{\prime}(a(\mu))$. Note that
$\lVert J^{\prime}(a(\mu))\bar{\times}_{3}(x-y)\rVert\leq\lVert J^{\prime}(a(\mu))\bar{\times}_{3}(x-y)\rVert_{\text{F}}=\lVert\mathtt{\mathbf{J}}(x-y)\rVert$
and $\lVert\mathtt{\mathbf{J}}(x-y)\rVert\leq\lVert\mathtt{\mathbf{J}}\rVert_{\text{F}}\lVert x-y\rVert=\lVert J^{\prime}(a(\mu))\rVert_{\text{F}}\lVert x-y\rVert$. Subsequently,
$\left\lVert J(x)-J(y)\right\rVert$ $=\left\lVert\int_{0}^{1}\,J^{\prime}(a(\mu))\bar{\times}_{3}(x-y)\,d\mu\right\rVert$
$\leq\int_{0}^{1}\,\left\lVert J^{\prime}(a(\mu))\right\rVert_{\text{F}}\left\lVert x-y\right\rVert\,d\mu$
$\leq\xi\left\lVert x-y\right\rVert,$
where $\xi=\max_{\mu\in[0,1]}\lVert J^{\prime}(a(\mu))\rVert_{\text{F}}<\infty$. This satisfies the Lipschitz condition. ∎
Corollary 5.1.2.
@@
Next, $\text{ker}\,(B)=\text{ker}\,(TB^{\dagger}KB)$ is tackled. If $x\in\text{ker}\,(B)$, then $Bx=0$. Thus, $TB^{\dagger}KBx=0$, and $x\in\text{ker}\,(TB^{\dagger}KB)$. Conversely, if $x\in\text{ker}\,(TB^{\dagger}KB)$, then $TB^{\dagger}KBx=0$. But $T^{-1}TB^{\dagger}KBx=B^{\dagger}KBx=0$. Since $\text{ker}\,(B^{\dagger})=\text{ker}\,(B^{\top})$, it suffices to consider $B^{\top}KBx=0$. Take $\kappa$ to be the smallest of all the (positive) diagonal entries of $K$. Subsequently, $\kappa(Bx)^{\top}(Bx)\leq(Bx)^{\top}K(Bx)=0$. Thus, $Bx=0$, and $x\in\text{ker}\,(B)$. By Proposition 2.1, the Jacobian is of rank n1n-1. ∎
@@
    Thus, $y^{\top}Hy>0$. ∎
A necessary condition (refer to [27, Theorem 2.2 on p. 15] and [28, Theorem 1.3.1 on p. 5]) for a local minimiser $x_{t}$ of $\theta_{c}$ is
@@
$x_{t}=x_{c}-\left(J(x_{c})^{\top}J(x_{c})+\lambda_{c}I\right)^{-1}J(x_{c})^{\top}g(x_{c}).$ (8)
$x_{t}$ is treated as a trial solution that could be the next iterate $x_{+}$ depending on how well the quadratic model approximates $\psi$. This is assessed through a comparison of the actual reduction
@@
If $\varrho=1$, then the quadratic model faithfully reproduces the behaviour of $\psi(x)$ around $x_{c}$. The trial solution is accepted for a value of $\varrho$ sufficiently greater than zero, and $\lambda_{c}$ is possibly decreased. Otherwise, $\lambda_{c}$ is increased and a new trial solution is computed. An algorithm detailing the steps in each iteration of the Levenberg-Marquardt method can be found in [28, Algorithm 3.3.5 on p. 58]. Quadratic convergence can be achieved [28, Theorem 3.3.4 on p. 58]. However, this is contingent on the sequence $\{x_{k}\}$ having the limit $x^{*}$ and $J(x^{*})$ being full rank.
Remark 5.
@@ 5.2 Rank deficiency and the Jacobian
The issue of a rank-deficient Jacobian in nonlinear least squares problems was examined in [29] but for non-zero (though small) $\psi(x^{*})$. Subset selection applied to the Jacobian was recommended over a truncated singular value decomposition. This involves the formation of a full rank variant, say $\widetilde{J}$, comprised solely of linearly independent columns from $J$. Instead of relying on (8), the system of linear equations
@@
There is the outstanding question though of how to discover the linearly independent columns of the Jacobian in the first place. A rank revealing algorithm could be employed for this task, but there is no need. After all, the underlying graph-based nature of the Nonlinear PageRank problem can be leveraged to determine the column in the Jacobian (see Corollary 5.1.2) that will play no part in the full rank variant. Since the problem solution figures in the graph partitioning process, it seems most prudent to fix the value of $x(v)$ for $v\in\mathcal{V}$ that is the furthest (minimum weighted) distance from the starting vertex $s\in\mathcal{V}$. Clearly, the relevant column in the Jacobian would then be the one omitted from $\widetilde{J}$.
Remark 6.
An alternative strategy could be to exclude the column in the Jacobian corresponding to the smallest entry of the minimal least squares solution to the Nonlinear PageRank problem at $p=2$ (see Proposition 4.3).